OpenStudy (anonymous):

If h is directly proportional to x^2, but inversely proportional to y, and h= 9/8 when x=1/2 and y=1/3, what is the value of h when x=1/3 and y=1/2?

OpenStudy (anonymous):

@Directrix @zzr0ck3r

OpenStudy (zzr0ck3r):

hmm

OpenStudy (anonymous):

I tried it, and got it wrong

OpenStudy (zzr0ck3r):

h = kx^2/y (9/8) = k* (1/4)*3 k = (9/8)*(4/3) = 3/2 so h = (3/2)x^2/y x=1/3 and y=1/2 h = (3/2)(1/9)*2 = 4/3 is that a choice?

Directrix (directrix):

@sakigirl

Directrix (directrix):

@sakigirl --> This is for you to work. See attachment. http://tinyurl.com/kbcny58

Directrix (directrix):

I posted a problem for @sakigirl to do - a bona fide SAT practice problem when she returns with the correct answer to the previous problem.

OpenStudy (anonymous):

c

Directrix (directrix):

@sakigirl While you were sleeping, you were asked a question: h = kx^2/y (9/8) = k* (1/4)*3 k = (9/8)*(4/3) = 3/2 so h = (3/2)x^2/y x=1/3 and y=1/2 h = (3/2)(1/9)*2 = 4/3 is that a choice? That is from @zzr0ck3r

Directrix (directrix):

So, is 4/3 right or not?

OpenStudy (anonymous):

@zzr0ck3r No , it isn't :(

OpenStudy (zzr0ck3r):

hmm im not sure what the question is asking. I am not sure if they want the same proportional constant